# A kinetic theory approach to consensus formation in financial markets

### v1

@@ 2.3.2. Short-run relationship with $D\ln(s(t))$ as the dependent variable
Table 3. Estimates of the short-term equation for the response variable $D\ln(s(t))$
***, **, *, $\cdot$ indicate stastistical significance at $0.1\%$, $1\%$, $5\%$, $10\%$ levels respectively
This equation is the most relevant one, as it demonstrates that the consensus target price for the S&\&&P 500 adjusts based on the market level, rather than the other way around. As a result, the consensus acts more as a “market follower” than a “market mover” in the short term.
@@
One can remark that $\alpha$ in the long-run equation is near 1. The Augmented Dickey-Fuller test for $\ln(s(t))-\ln(X(t))$ leads to the conclusion that one can accept that $\alpha$ is equal to 1, which, again, is no surprise when looking at Figure 1.
@@
All these results permits us to design our kinetic model which, in return, will improve our understanding of the long-term dynamics linking the consensus to the market.
@@ 3. The kinetic model
The kinetic-type model described in this Section predicts the behavior of the S&P 500 index by supposing that the influence on the index forecasts is governed by a small number of inputs, namely the opinions of the other agents (consensus), the market suggestions, and the degree of attractiveness of the share price.
@@
The agents’ community is described by the distribution function $f=f(t,x)$, defined on +×$\mathbb{R}^{+}\times\bar{\Omega}$. This function represents the density of agents holding price forecast $x$ at time $t$. Its evolution is governed by a kinetic equation, encompassing two levels of interaction: the microscopic level, where opinions on the price forecasts are influenced by the mean field and external market forces, and the macroscopic level, which captures the evolution of $f(t,x)$ induced by these microscopic interactions.
@@
    Attractiveness. The attractiveness is described by a parameter $q\in(0,1)$ that quantifies the degree to which a one-year price forecast $x$ is drawn toward the share price $X(t)$.
@@
where $\Delta\in\mathbb{R}$ is a spread function that encapsulates the deviation of forecasts from the actual market price.
@@
We first introduce the sentiment of the population on the target price $s\,:\,\mathbb{R}^{*}_{+}\to\mathbb{R}_{+}$, defined as:
@@
(5) $\left\{\begin{array}[]{l}\displaystyle\partial_{t}f+\alpha\partial_{x}\big{(}(s(t)-x)f\big{)}=\displaystyle\beta\left[\frac{1}{1-q}f\left(\frac{x}{1-q}-\frac{q}{1-q}X(t)(1+\Delta)\right)\mathds{1}_{\left(\frac{x}{1-q}-\frac{q}{1-q}X(t)(1+\Delta)\geq 0\right)}-f(x)\right]\\[10.0pt] f(0,x)=f^{\text{in}}(x)\qquad x\in{\mathbb{R}^{*}_{+}}\\[10.0pt] f(t,0)=0\qquad t\in{\mathbb{R}^{*}_{+}},\\[10.0pt] \end{array}\right.$
where $\alpha\in\mathbb{R}_{+}^{*}$ is a parameter which describes how strong is the comparison with the sentiment of the population and $\beta\in\mathbb{R}_{+}^{*}$ is the characteristic relaxation time to the spot price.
The weak formulation
It is convenient to express and to study the previous model in weak form. Let $\phi\in C_{c}^{1}(\mathbb{R}\times\mathbb{R})$, which is the space of funtions of class $C^{1}$ with compact support in $\mathbb{R}\times\mathbb{R}$. From (5), we deduce formally that
(6) $\begin{array}[]{l}\displaystyle\frac{\mathrm{d}}{{\rm d}t}\int_{\mathbb{R}_{+}}f(t,x)\phi(t,x){\rm d}x+\alpha\int_{\mathbb{R}_{+}}\big{(}(s(t)-x)f\big{)}\partial_{x}\phi(t,x){\rm d}x=\displaystyle\beta\int_{\mathbb{R}_{+}}f(t,x)\left[\phi(t,\bar{x})-\phi(t,x)\right]{\rm d}x\end{array}$
for all $\phi\in C_{c}^{1}(\mathbb{R}\times\mathbb{R})$.
@@
When $\phi=1$, we deduce from the weak form that
$\frac{\mathrm{d}}{{\rm d}t}m_{0}=0,$
@@
On the other hand, when $\phi=x$, the weak form (6) reduces to the following ordinary differential equation governing the time evolution of $m_{1}$:
@@
If we divide by the constant $m_{0}$, we deduce an ordinary differential equation which describes the time evolution of the sentiment of the population:
@@ 4. Existence and uniqueness of the solution
$\partial_{t}f+\alpha\left(s(t)-x\right)\partial_{x}f=\frac{\beta}{1-q}f\left(\frac{x}{1-q}-\frac{q}{1-q}X(t)(1+\Delta)\right)\mathds{1}_{\left(\frac{x}{1-q}-\frac{q}{1-q}X(t)(1+\Delta)\geq 0\right)}+(\alpha-\beta)f(t,x).$
@@
(8) $\begin{array}[]{l}\displaystyle f(t,x)=f^{\text{in}}\left(xe^{\alpha t}-\alpha\int_{0}^{t}s(\theta)e^{\alpha\theta}\mathrm{d}\theta\right)e^{(\alpha-\beta)t}\mathds{1}_{\left[\left(xe^{\alpha t}-\alpha\int_{0}^{t}s(\theta)e^{\alpha\theta}\mathrm{d}\theta\right)\geq 0\right]}+\\[10.0pt] \displaystyle\frac{\beta}{1-q}\int_{0}^{t}e^{(\beta-\alpha)(\theta-t)}f\left(\frac{x}{1-q}-\frac{q}{1-q}X(\theta)(1+\Delta)\right)\mathds{1}_{\left(\frac{x}{1-q}-\frac{q}{1-q}X(\theta)(1+\Delta)\geq 0\right)}\mathrm{d}\theta,\end{array}$
@@
Let $T\in\mathbb{R}_{+}^{*}$ and $C_{b}([0,T]\times\mathbb{R}_{+})$ the space of uniformly bounded continuous functions on $[0,T]\times\mathbb{R}_{+}$, such that
@@
$F(f^{\text{in}})=f^{\text{in}}\left(xe^{\alpha t}-\alpha\int_{0}^{t}s(\theta)e^{\alpha\theta}\mathrm{d}\theta\right)e^{(\alpha-\beta)t}\mathds{1}_{\left[\left(xe^{\alpha t}-\alpha\int_{0}^{t}s(\theta)e^{\alpha\theta}\mathrm{d}\theta\right)\geq 0\right]}$
@@
$\mathscr{T}g=\frac{\beta}{1-q}\int_{0}^{t}e^{(\beta-\alpha)(\theta-t)}g\left(\frac{x}{1-q}-\frac{q}{1-q}X(\theta)(1+\Delta)\right)\mathds{1}_{\left(\frac{x}{1-q}-\frac{q}{1-q}X(\theta)(1+\Delta)\geq 0\right)}\mathrm{d}\theta,$
@@
$\begin{array}[]{c}\displaystyle\|f\|_{C_{b}([0,T]\times\mathbb{R}_{+})}\leq\displaystyle\sum_{n=0}^{+\infty}\|\mathscr{T}^{n}\|_{\mathcal{L}(C_{b}([0,T]\times\mathbb{R}_{+}))}\|F(f^{\text{in}})\|_{C_{b}([0,T]\times\mathbb{R}_{+})}\leq\\[10.0pt] \displaystyle\sum_{n=0}^{+\infty}\|f^{\text{in}}\|_{C_{b}([0,T]\times\mathbb{R}_{+})}e^{|\alpha-\beta|T}\left(\frac{\beta T}{1-q}\right)^{n}\frac{1}{n!}=\|f^{\text{in}}\|_{C_{b}([0,T]\times\mathbb{R}_{+})}e^{|\alpha-\beta|T}\exp\left(\frac{\beta}{1-q}T\right).\\ \end{array}$
@@
$\|\tilde{f}\|_{C_{b}([0,T]\times\mathbb{R}_{+})}=0\text{ or }\|\mathscr{T}^{n}\tilde{f}\|_{C_{b}([0,T]\times\mathbb{R}_{+})}$
@@ 5.1. Time evolution of the price sentiment
(11) $\left\{\begin{array}[]{l}\displaystyle\frac{\mathrm{d}}{{\rm d}t}s(t)=q\beta\big{[}X(t)(1+\Delta)-s(t)\big{]}\\[10.0pt] \displaystyle s(0)=s^{\text{in}}=\left(\int_{\mathbb{R}_{+}}xf^{\text{in}}(x)\mathrm{d}x\right)\\ \end{array}\right.$
@@
$s(t)=\left(\int_{\mathbb{R}_{+}}xf^{\text{in}}(x)\mathrm{d}x\right)\left(\int_{\mathbb{R}_{+}}f^{\text{in}}(x)\mathrm{d}x\right)^{-1}\ e^{-q\beta t}+q\beta\int_{0}^{t}X(\theta)(1+\Delta)e^{q\beta(\theta-t)}\mathrm{d}\theta.$
We have developed a MATLAB script that implements a numerical solution for the differential equation (11) using a standard fourth-order Runge-Kutta method. We recall that the spot price $X(t)$ is derived from time-series data of the S&P 500 index, covering the period from May 26th, 2009, to November 20th, 2024.
To calibrate the model parameters and assess its accuracy, we divided the dataset into two subsets. The first subset, referred to as the training set, includes S&P 500 index data from May 26, 2009, to December 31, 2016. For the sentiment data $s_{\text{meas}}(t)$, we used the median of one-year analysts’ forecasts for the S&P 500 index over the same period. The training set consists of 1,987 data points.
@@
Using the optimized parameters, we simulated the period from January 3rd, 2017, to November 20th, 2024, and compared the results with the corresponding time series, consisting of 2,055 data points.
Figure 3 presents the results of our model compared to the median of one-year analysts’ forecasts for the forecasting set, which spans from January 3, 2017, to November 20, 2024, and to the cointegration model equilibrium which parameters are estimated over the same period as the kinetic model. For this simulation, the sentiment $s(0)$ was initialized using the value computed on January 3, 2017, $s(0)=2463.93$. It is clear that the kinetic model better captures the long-term dynamics compared to the cointegration model. Moreover, while the long-term equilibrium estimated by the cointegration model is much more volatile, closely mirroring the fluctuations of the S&\&&P 500 (due to $\alpha$ being near 1), the kinetic model does not suffer from the same volatility issues.
Figure 3. Forecasts $s_{\text{meas}}$ of the kinetic model (gray line), Forecasts $\exp{(\alpha\ln(X(t))+\beta)}$ of the cointegration model (dark blue line) and the time series of the median of one-year analysts’ forecasts for the S&P 500 index $s(t)$ (light blue line), from January 3rd, 2017, to November 20th, 2024.
Finally, in Figure 4, we plot the relative error between the forecasts generated by our model and the real data for the median of one-year analysts’ forecasts.
The statistical analysis of the relative errors reveals several insights into the robustness of the forecast. The mean relative error is equal to -0.00168, while the median is equal to -0.00217, indicating a slight systematic underestimation. This bias can be the consequence of the different market trends of the training set with respect to the forecasting set. The standard deviation of the errors, 0.00111, reflects a low level of dispersion, showing that the predictions are consistently close to the actual values.
@@
Figure 4. Analysis of the relative error between the forecasts $s(t)$ of the model and the time series of the median of one-year analysts’ forecasts for the S&P 500 index $s_{\text{meas}}$, from January 3rd, 2017, to November 20th, 2024.
Acknowledgements. FS acknowledges the support of INdAM, GNFM group.
Conflicts of interest. The authors state that there is no conflict of interest.

### v2

@@ 3.2.3. Short-run relationship with $D\ln(s_{\text{meas}}(t))$ as the dependent variable
This equation is the most relevant one, as it demonstrates that the consensus target price for the S&\&P 500 adjusts based on the market level, rather than the other way around. As a result, the consensus (i.e. the one-year target price) acts more as a “market follower” than a “market mover” in the short term.
@@
One can remark that $\alpha$ in the long-run equation is near 1. The Augmented Dickey-Fuller test for $\ln(s_{\text{meas}}(t))-\ln(X(t))$ leads to the conclusion that one can accept that $\alpha$ is equal to 1, which, again, is no surprise when looking at Figure 1.
@@
3.3. Diagnostic tests for residuals
To perform the diagnostic tests, we first need to convert the VECM (Vector Error Correction Model) into its CVAR (Cointegrated Vector Autoregression) form. We use the vec2var procedure in R to complete this transformation and retrieve the necessary residuals
3.3.1. Serial Correlation
The Breusch-Godfrey serial correlation test of the residuals (see table 10 below) confirms the absence of autocorrelation, validating that the lag length in the model was correctly specified.
Table 10. Breusch-Godfrey test
Statistic p-value
49.636 0.14
3.3.2. Check for normality
In contrast, the normality assumption for residuals is rejected (see table 11). This is a known result, at least regarding the log returns of the S&\&P 500. Indeed, even though this assumption is implicit in most financial models (Markowitz Optimization, Black-Scholes Equation), it is generally not met for stock returns. Nevertheless, given our large sample size, the Central Limit Theorem is relied upon to ensure the validity of our inferences.
Table 11. Jarque-Bera test
Statistic p-value
76932 0.0000
3.3.3. Arch effect
Conversely, heteroscedasticity in the residuals is observed (see Table 12). This issue is likely attributable to the volatility regime changes in equity markets, the same phenomenon that necessitated retaining a large number of lags in the model. Ideally, the equilibrium model should be estimated on a rolling basis and re-estimated regularly. However, even though this procedure improves the test statistics for ARCH effect, the homoscedasticity assumption is consistently rejected, regardless of a reduced sample size. This persistence is arguably the model’s main weakness. Despite this limitation, the approach proved valuable for understanding the underlying dynamics of consensus formation, allowing us to successfully calibrate the kinetic model (see next section), which inherently requires no assumptions about the distribution of log returns or homoscedasticity.
Table 12. Heteroscedasticity test for residuals
Statistic p-value
1613 0.0000
@@ 4. The kinetic model
The kinetic-type model described in this Section predicts the behavior of the analysts’ recommendations on the one-year S&P 500 index by supposing that the influence on the index forecasts is governed by a small number of inputs, namely the opinions of the other agents (consensus), the market suggestions, and the degree of attractiveness of the share price.
@@
The agents’ community is described by the distribution function $f=f(t,x)$, defined on $\mathbb{R}^{+}\times\bar{\Omega}$. This function represents the density of agents holding price forecast $x$ at time $t$. Its evolution is governed by a kinetic equation, encompassing two levels of interaction: the microscopic level, where opinions on the price forecasts are influenced by the mean field and external market forces, and the macroscopic level, which captures the evolution of $f(t,x)$ induced by these microscopic interactions.
@@
    Attractiveness. The attractiveness is described by a parameter $q\in(0,1)$ that quantifies the degree to which a one-year price forecast $x$ is drawn toward the share price $X(t)$ augmented by a premium.
@@
where $\Delta\in\mathbb{R}$ is the spread that encapsulates the deviation of forecasts from the actual market price.
@@
We first introduce the function $s\,:\,\mathbb{R}^{*}_{+}\to\mathbb{R}_{+}$, which describes the average recommendation of the analysts’ population on the target price, defined as:
@@
In what follows, this quantity will sometimes be indicated as “the sentiment”.
@@
[…]
where $\alpha\in\mathbb{R}_{+}^{*}$ is a parameter which describes how strong is the comparison with $s$ and $\beta\in\mathbb{R}_{+}^{*}$ is the characteristic relaxation time to the spot price.
The semi-weak formulation
It is convenient to express and to study the previous model in semi-weak form. Let $\phi\in C^{1}(\mathbb{R})$ and suppose that
$\lim_{x\to 0^{+}}\phi(x)f(t,x)=\lim_{x\to+\infty}\phi(x)f(t,x)=0\text{ for all }t\in\mathbb{R}_{+}\text{ and for all }\phi\in{C^{1}(\mathbb{R})}.$
The previous assumption implies that no analyst assigns a one-year target price for the S&P 500 index equal to zero or infinity.
From (5), we deduce formally that
(6) $\begin{array}[]{l}\displaystyle\frac{\mathrm{d}}{{\rm d}t}\int_{\mathbb{R}_{+}}f(t,x)\phi(x){\rm d}x+\alpha\int_{\mathbb{R}_{+}}\big((s(t)-x)f\big)\partial_{x}\phi(x){\rm d}x=\displaystyle\beta\int_{\mathbb{R}_{+}}f(t,x)\left[\phi(\bar{x})-\phi(x)\right]{\rm d}x\end{array}$
for all $\phi\in{C^{1}(\mathbb{R})}$.
@@
When $\phi=1$, we deduce from the semi-weak form that
$\frac{\mathrm{d}}{{\rm d}t}m_{0}(t)=0,$
@@
On the other hand, when $\phi=x$, the semi-weak form (6) reduces to the following ordinary differential equation governing the time evolution of $m_{1}$:
@@
If we divide by the constant $m_{0}$, we deduce an ordinary differential equation which describes the time evolution of the sentiment of the analysts’ population:
@@
Variance and risk implications
A key advantage of the kinetic model is that it does not only provide the mean behavior of the system, but it also encapsulates richer distributional information. For instance, it allows us to compute higher-order moments such as the variance, as detailed below, which are crucial for understanding uncertainty and predicting risk or volatility.
In order to compute the variance, we begin with introducing the second-order moment
$m_{2}(t)=\int_{\mathbb{R}_{+}}x^{2}f(t,x)\,\mathrm{d}x.$
By definition, the variance is given by
(8) $\mathcal{V}(t)=\frac{m_{2}(t)}{m_{0}}-\left(\frac{m_{1}(t)}{m_{0}}\right)^{2},$
where $m_{0}$ and $m_{1}$ are the zeroth and first-order moments previously defined.
Applying the weak formulation (6) with the test function $\phi(x)=x^{2}$, we obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}_{+}}x^{2}f(t,x)\,\mathrm{d}x+\alpha\int_{\mathbb{R}_{+}}(s(t)-x)f(t,x)\,(2x)\,\mathrm{d}x=\beta\int_{\mathbb{R}_{+}}f(t,x)\,\left(\bar{x}^{2}-x^{2}\right)\,\mathrm{d}x,$
i.e.
$\frac{\mathrm{d}}{\mathrm{d}t}m_{2}(t)+2\alpha\big[s(t)m_{1}(t)-m_{2}(t)\big])=\beta\left[(q^{2}-2q)\,m_{2}(t)+2q(1-q)X(t)(1+\Delta)m_{1}(t)+q^{2}X(t)^{2}(1+\Delta)^{2}m_{0}\right],$
where we used that the quantities $q$, $X$ and $\Delta$ are independent of $x$.
Since $s(t)=m_{1}(t)/m_{0}$, we rewrite the previous equation as
(9) $\frac{\mathrm{d}}{\mathrm{d}t}m_{2}(t)=\left[2\alpha+\beta(q^{2}-2q)\right]m_{2}(t)-2\alpha\frac{m_{1}(t)^{2}}{m_{0}}+2\beta{q}(1-q)X(t)(1+\Delta)m_{1}(t)+\beta q^{2}X(t)^{2}(1+\Delta)^{2}m_{0}.$
Combining (8) and (9), using the ODE for $m_{1}$ and the definition of $s$, given in (4), we deduce that $\mathcal{V}$ satisfies
(10) $\frac{\mathrm{d}}{\mathrm{d}t}\,\mathcal{V}(t)=\left[2\alpha+\beta(q^{2}-2q)\right]\mathcal{V}(t)+\beta(q^{2}{-2q})s(t)^{2}+2\beta{q}(1-2q)\,X(t)(1+\Delta)s(t)+\beta q^{2}\,X(t)^{2}(1+\Delta)^{2}$
with initial condition
$\mathcal{V}(0)=\frac{m_{2}^{\mathrm{in}}}{m_{0}}-\left(\frac{m_{1}^{\mathrm{in}}}{m_{0}}\right)^{2},$
where
$m_{1}^{\mathrm{in}}=\int_{\mathbb{R}_{+}}xf^{\mathrm{in}}(x){\rm d}x\text{\quad and\quad}m_{2}^{\mathrm{in}}=\int_{\mathbb{R}_{+}}x^{2}f^{\mathrm{in}}(x){\rm d}x.$
The solution of (10) is
(11) $\mathcal{V}(t)=e^{\gamma t}\left[\mathcal{V}(0)+\int_{0}^{t}e^{-\gamma\tau}\left(\beta(q^{2}{-2q})s(\tau)^{2}+2\beta{q}(1-2q)\,X(\tau)(1+\Delta)s(\tau)+\beta q^{2}\,X(\tau)^{2}(1+\Delta)^{2}\right)\,\mathrm{d}\tau\right],$
where $s$ is given by (7) and
$\gamma=2\alpha+\beta(q^{2}-2q).$
We immediately see that the sign of the coefficient $\gamma$ determines the intrinsic stability of the variance dynamics. In particular, when $\gamma<0$, the linear term in (10) acts as a damping factor. If the forcing term vanishes (i.e., the integral term), the variance decays exponentially. This corresponds to a mean-reverting behavior where dispersion and risk tend to stabilize over time.
On the other hand, if $\gamma>0$, the regime is unstable. When $2\alpha>\beta q(2-q)$, the internal interaction dominates the consensus mechanism, and the variance grows exponentially even in the absence of external shocks. This scenario reflects an intrinsically unstable system, where disagreement amplifies over time.
Therefore, the kinetic model does not only provide the mean sentiment (i.e. the average analysts’ recommendation) but also its distributional spread, which is a key forward-looking indicator of risk and volatility. In financial applications, such dispersion-based measures have been shown to anticipate periods of elevated uncertainty [7], making $\mathcal{V}(t)$ a natural proxy for volatility forecasting and risk management.
@@ 5. Existence and uniqueness of the solution
$\partial_{t}f+\alpha\left(s(t)-x\right)\partial_{x}f=\frac{\beta}{1-q}f\left(\frac{x}{1-q}-\frac{q}{1-q}X(t)(1+\Delta)\right)\mathbf{1}_{\left(\frac{x}{1-q}-\frac{q}{1-q}X(t)(1+\Delta)\geq 0\right)}+(\alpha-\beta)f(t,x).$
@@
(12) $\begin{array}[]{l}\displaystyle f(t,x)=f^{\text{in}}\left(xe^{\alpha t}-\alpha\int_{0}^{t}s(\theta)e^{\alpha\theta}\mathrm{d}\theta\right)e^{(\alpha-\beta)t}\mathbf{1}_{\left[\left(xe^{\alpha t}-\alpha\int_{0}^{t}s(\theta)e^{\alpha\theta}\mathrm{d}\theta\right)\geq 0\right]}+\\[10.0pt] \displaystyle\frac{\beta}{1-q}\int_{0}^{t}e^{(\beta-\alpha)(\theta-t)}f\left(\frac{x}{1-q}-\frac{q}{1-q}X(\theta)(1+\Delta)\right)\mathbf{1}_{\left(\frac{x}{1-q}-\frac{q}{1-q}X(\theta)(1+\Delta)\geq 0\right)}\mathrm{d}\theta,\end{array}$
@@
Let $T\in\mathbb{R}_{+}^{*}$ and $C_{b}([0,T]\times\mathbb{R}_{+})$ the space of uniformly bounded continuous functions on $0,T]\times\mathbb{R}_{+}$, such that
@@
$F(f^{\text{in}})=f^{\text{in}}\left(xe^{\alpha t}-\alpha\int_{0}^{t}s(\theta)e^{\alpha\theta}\mathrm{d}\theta\right)e^{(\alpha-\beta)t}\mathbf{1}_{\left[\left(xe^{\alpha t}-\alpha\int_{0}^{t}s(\theta)e^{\alpha\theta}\mathrm{d}\theta\right)\geq 0\right]}$
@@
$\mathscr{T}g=\frac{\beta}{1-q}\int_{0}^{t}e^{(\beta-\alpha)(\theta-t)}g\left(\frac{x}{1-q}-\frac{q}{1-q}X(\theta)(1+\Delta)\right)\mathbf{1}_{\left(\frac{x}{1-q}-\frac{q}{1-q}X(\theta)(1+\Delta)\geq 0\right)}\mathrm{d}\theta,$
@@
[…]
@@
$\|\tilde{f}\|_{C_{b}([0,T]\times\mathbb{R}_{+})}=0\text{ or }\|\mathscr{T}^{n}\tilde{f}\|_{C_{b}([0,T]\times\mathbb{R}_{+})}.$
@@ 6.1. Time evolution of the price sentiment
[…]
@@
$s(t)=\left(\int_{\mathbb{R}_{+}}xf^{\text{in}}(x)\mathrm{d}x\right)\left(\int_{\mathbb{R}_{+}}f^{\text{in}}(x)\mathrm{d}x\right)^{-1}\ e^{-q\beta t}+q\beta\int_{0}^{t}X(\theta)(1+\Delta)e^{q\beta(\theta-t)}\mathrm{d}\theta$
We have developed a MATLAB script that implements a numerical solution for the differential equation (15) using a standard fourth-order Runge-Kutta method. We recall that the spot price $X(t)$ is derived from time-series data of the S&P 500 index, covering the period from May 26th, 2009, to June 2nd, 2025.
To calibrate the model parameters and assess its accuracy, we divided the dataset into two subsets. The first subset, referred to as the training set, includes S&P 500 index data from May 26th, 2009, to December 31st, 2016. For the sentiment data $s_{\text{meas}}(t)$, we used the average of one-year analysts’ forecasts for the S&P 500 index over the same period. The training set consists of 1,987 data points.
@@
Using the optimized parameters, we simulated the period from January 3rd, 2017, to June 2nd, 2025, and compared the results with the corresponding time series, consisting of 2,193 data points.
Figure 3 presents the results of our model compared to the average of one-year analysts’ forecasts for the forecasting set, which spans from January 3rd, 2017, to June 2nd, 2025, and to the cointegration model equilibrium which parameters are estimated over the same period as the kinetic model. For this simulation, the sentiment $s(0)$ was initialized using the value computed on January 3, 2017, $s(0)=2463.93$. It is clear that the kinetic model better captures the long-term dynamics compared to the cointegration model. Moreover, while the long-term equilibrium estimated by the cointegration model is much more volatile, closely mirroring the fluctuations of the S&\&P 500 (due to $\alpha$ being near 1), the kinetic model does not suffer from the same volatility issues.
Figure 3. Forecasts $s(t)$ of the kinetic model (black line), Forecasts $\exp{(\alpha\ln(X(t))+\beta)}$ of the cointegration model (grey line) and the time series of the average of one-year analysts’ forecasts for the S&P 500 index $s_{\text{meas}}(t)$ (light grey line), from January 3rd, 2017, to June 2nd, 2025.
Finally, in Figure 4, we plot the relative error between the forecasts generated by our model and the real data for the average of one-year analysts’ forecasts.
Figure 4. Analysis of the relative error between the forecasts $s(t)$ of the model and the time series of the average of one-year analysts’ forecasts for the S&P 500 index $s_{\text{meas}}$, from January 3rd, 2017, to June 2nd, 2025.
The statistical analysis of the relative errors reveals several insights into the robustness of the forecast. The mean relative error is equal to 0,015485, while the median is equal to 0,014096, indicating a slight systematic overestimation. This bias can be the consequence of the different market trends of the training set with respect to the forecasting set. The standard error, 0,000213, reflects a low level of dispersion, showing that the predictions are consistently close to the actual values.
@@
7. Concluding remarks
[…]
This initial intuition was informed by the VECM analysis, but it is the kinetic model that allows for a precise description of how analysts weigh the consensus opinion relative to the current price. Crucially, it appears analysts place greater weight on the market price than on the opinions of their peers.
This comparative analysis decisively establishes the kinetic model’s superiority over the traditional VECM for this specific forecasting task. The VECM is severely limited by the necessity of restrictive assumptions like normality and homoscedasticity – assumptions frequently violated in volatile equity data – which cause it to suffer from significant forecasting deviations.
In sharp contrast, the kinetic model is inherently more robust: its design specifically targets out-of-equilibrium phenomena, requiring no underlying distributional hypotheses. Consequently, the kinetic model provides highly stable and convincing forecasts, maintaining its efficacy even during chaotic and unpredictable market periods and over a long-term horizon.
Use of AI tools declaration
The English in this paper has been refined using Artificial Intelligence tools.
Acknowledgments
FS acknowledges the support of INdAM, GNFM group.
Conflict of interest
The authors state that there are no conflicts of interest nor ethical issues.